You short sold 1,200 shares of stock at a price of $30 and an initial margin of 75 percent. If the maintenance margin is 30 percent, at what share price will you receive a margin call? What is your account equity at this stock price? (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Margin call price
Account equity

2 decimal places required.
Answer #1

Proceeds from short sale = 1,200×$30 = $36,000
Initial deposit = $36,000×0.75 = $27,000
Account value = $36,000 + 27,000 = $63,000
Margin call price = $63,000/[1,200+(0.3 × 1,200)] = $40.38
Account equity = $63,000 - (1,200 × $40.38) = $14,544

Hence, Margin call price is $40.38 and Account equity is $14,544.